OpenStudy (anonymous):

A student club has 20 members. How many ways can the club choose a President, Secretary, and Treasurer? A. 6,840 B. 8,000 C. 2,280

OpenStudy (anonymous):

Would you have to make a table? (:

OpenStudy (anonymous):

no just mulitply \[20\times 19\times 18\]

OpenStudy (anonymous):

Ahhhh ok :)

OpenStudy (anonymous):

P(8, 5) = A. 40,320 B. 1,680 C. 6,720 so then wouldnt the answer to this one be A?

OpenStudy (anonymous):

This is a combination/permutation problem. Since "order" does matter, ie, if Sally, Bob and Mike are the three officers, there can be multiple ways to fill the positions: Sally could be Pres, Bob Sec, and Mike Treas, or Sally could be Sec, Bob Pres, Mike Treas, and so on. Thus, this is a permutation. If order did not matter, ie, all three held the same title, then it would be a combination. A permutation can be written as nPr where n=total group, P means permutation, and r=number chosen. Thus, this would be 20P3.

OpenStudy (anonymous):

Ohhh >.< so satellite was wrong?

OpenStudy (anonymous):

satellite is correct, but I am giving you the long answer. :P

OpenStudy (anonymous):

Ahhh yea ok :) So for the second one wouldnt it be A? (:

OpenStudy (anonymous):

Permutation is calculated as n!/(n-r)! For #1, this is 20!/(20-3)! = 20*19*18. For #2, it is 8!/(8-5)!. This equals 8*7*6*5*4*3*2*1/(3*2*1) which equals 8*7*6*5*4. See why? So A is not correct...

OpenStudy (anonymous):

Do you know what is?

OpenStudy (anonymous):

Yes it would be C correct?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Woo!! 100%! :) Thanks r u a teacher? :D

OpenStudy (anonymous):

yw. No, not a teacher.

OpenStudy (anonymous):

Well you definitely should be one! (: Both of you!

OpenStudy (anonymous):

Thanks. ;)

OpenStudy (anonymous):

Yea getting ready to take a assessment on my online school so i might be posting a few more questions ^_^

OpenStudy (anonymous):

cheater! lol

OpenStudy (anonymous):

Haha not cheating if ur old computer hard drive crashed and now u cant access ur notes!! >.< and finals are almost here too! im soooo stressed all my notes. just, gon.

OpenStudy (anonymous):

gone*